II. EXERCISE: Directions: Solve the problems and show your complete solution.

1. What volume of 23% m/v solution contains 453 g of solute?

2. Find the amount of solute to make 0.5 L of a 39% m/v solution.


1
Expert's answer
2021-09-17T02:13:04-0400

1. 453 = 100%

? = 23%

= 453 × 23/100 = 104.19ml


2. 0.5 L = 39%

500ml = 39%

? = 100%

= 500×100/39

= 1282.1g